Write the slope-intercept equation for the line that passes through (2, -7) and (4, -13). Please show all of your work.
Answer by rfer(16322)

m= (-13--7)/(4-2)
m=-6/2
m=-3
---------------
y=mx+b
-7=-3*2+b
b=-1
----------------
y=-3x-1